# Resonance Contributions to Charged-Current (Anti)Neutrino-Nucleon Elastic Scattering at GeV Energies

### v1

@@ 5 Conclusions and outlook
In this study, we evaluated the $\Delta(1232)$-resonance contribution to the charged-current (anti)neutrino-nucleon elastic scattering for the first time at accelerator neutrino energies. We modeled the $\Delta$ interaction vertices in the on-shell form, with and without shifts in momenta for the electroweak interaction vertex. We approximated the vector part of the nucleon-to-delta transition current by the leading magnetic dipole transition. We determined the virtual contributions of the $\Delta$ intermediate state to all invariant amplitudes. Throughout the calculation, we verified that the resulting amplitudes respected the expected soft, collinear, and crossing symmetry properties. We also ensured the consistency with unitarity conditions for the invariant amplitudes of inelastic excitations in the charged-current (anti)neutrino-nucleon elastic scattering. Our results show that the $\Delta$-resonance contribution remains at the permille level for the unpolarized elastic scattering cross sections, which is in agreement with the error estimates presented in previous works [83, 84]. The analysis further confirms no significant kinematic enhancements of the resonance contribution. Not only do the central values of radiative corrections are validated, but the corresponding uncertainties from the earlier studies are also confirmed, ensuring that the predictions made in Refs. [83, 84] remain the most precise and reliable for future experiments and studies.
@@ Appendix A Relation between helicity and invariant amplitudes
To describe charged-current (anti)neutrino-nucleon elastic scattering, there are $8$ helicity amplitudes $T_{\lambda^{\prime}h^{\prime},\lambda}$ with arbitrary helicities of the charged lepton, final and initial nucleons $h^{\prime},\lambda,\lambda^{\prime}$ = $\pm 1/2$, respectively: $T_{1}=T_{{\frac{1}{2}}{\frac{1}{2}},{\frac{1}{2}}},\penalty 10000\ T_{2}=T_{{\frac{1}{2}}{-\frac{1}{2}},{\frac{1}{2}}},\penalty 10000\ T_{3}=T_{{-\frac{1}{2}}{\frac{1}{2}},{-\frac{1}{2}}},\penalty 10000\ T_{4}=T_{{-\frac{1}{2}}{-\frac{1}{2}},{-\frac{1}{2}}},\penalty 10000\ T_{5}=T_{{\frac{1}{2}}{\frac{1}{2}},{-\frac{1}{2}}},\penalty 10000\ T_{6}=T_{{\frac{1}{2}}{-\frac{1}{2}},{-\frac{1}{2}}},\penalty 10000\ T_{7}=T_{{-\frac{1}{2}}{\frac{1}{2}},{\frac{1}{2}}},\penalty 10000\ T_{8}=T_{{-\frac{1}{2}}{-\frac{1}{2}},{\frac{1}{2}}}$.
Using the Jacob and Wick [148] phase convention for the spinors, the helicity amplitudes $T_{\lambda^{\prime}h^{\prime},\lambda}$ for (anti)neutrino-nucleon elastic scattering can be expressed in terms of the invariant amplitudes as
@@ Acknowledgements
Author thanks Barbara Pasquini, Marc Vanderhaeghen, and Lothar Tiator for numerous discussions while working with pion electroproduction amplitudes, Kevin Quirion and Emilie Passemar for numerous discussions while working with pion neutrinoproduction. This work is supported by the US Department of Energy through the Los Alamos National Laboratory and by LANL’s Laboratory Directed Research and Development (LDRD/PRD) program under project numbers 20210968PRD4 and 20240127ER. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the US Department of Energy (Contract No. 89233218CNA000001). FeynCalc [149, 150], LoopTools [151], Mathematica [152], and DataGraph [153] were extremely useful in this work.

### v3

@@ 6 Conclusions and outlook
In this study, we evaluated the $\Delta(1232)$-resonance contribution to the charged-current elastic (anti)neutrino-nucleon scattering for the first time at accelerator neutrino energies. We modeled the $\Delta$ interaction vertices in the on-shell form, with and without shifts in momenta for the electroweak interaction vertex. We approximated the vector part of the nucleon-to-delta transition current by the leading magnetic dipole transition. We determined the virtual contributions of the $\Delta$ intermediate state to all invariant amplitudes. Throughout the calculation, we verified that the resulting amplitudes respected the expected soft, collinear, and crossing symmetry properties. We also ensured the consistency with unitarity conditions for the invariant amplitudes of inelastic excitations in the charged-current elastic (anti)neutrino-nucleon scattering. Our results show that the $\Delta$-resonance contribution remains at the permille level for the unpolarized elastic scattering cross sections, which is in agreement with the error estimates presented in previous works [84, 85]. The analysis further confirms no significant kinematic enhancements of the resonance contribution. Not only do the central values of radiative corrections are validated, but the corresponding uncertainties from the earlier studies are also confirmed, ensuring that the predictions made in Refs. [84, 85] remain the most precise and reliable for modern and future experiments and studies.
@@ Appendix A Relation between helicity and invariant amplitudes
To describe charged-current elastic (anti)neutrino-nucleon scattering, there are $8$ helicity amplitudes $T_{\lambda^{\prime}h^{\prime},\lambda}$ with arbitrary helicities of the charged lepton, final and initial nucleons $h^{\prime},\lambda,\lambda^{\prime}$ = $\pm 1/2$, respectively: $T_{1}=T_{{\frac{1}{2}}{\frac{1}{2}},{\frac{1}{2}}},\penalty 10000\ T_{2}=T_{{\frac{1}{2}}{-\frac{1}{2}},{\frac{1}{2}}},\penalty 10000\ T_{3}=T_{{-\frac{1}{2}}{\frac{1}{2}},{-\frac{1}{2}}},\penalty 10000\ T_{4}=T_{{-\frac{1}{2}}{-\frac{1}{2}},{-\frac{1}{2}}},\penalty 10000\ T_{5}=T_{{\frac{1}{2}}{\frac{1}{2}},{-\frac{1}{2}}},\penalty 10000\ T_{6}=T_{{\frac{1}{2}}{-\frac{1}{2}},{-\frac{1}{2}}},\penalty 10000\ T_{7}=T_{{-\frac{1}{2}}{\frac{1}{2}},{\frac{1}{2}}},\penalty 10000\ T_{8}=T_{{-\frac{1}{2}}{-\frac{1}{2}},{\frac{1}{2}}}$.
Using the Jacob and Wick [151] phase convention for the spinors, the helicity amplitudes $T_{\lambda^{\prime}h^{\prime},\lambda}$ for elastic (anti)neutrino-nucleon scattering can be expressed in terms of the invariant amplitudes as
@@ Acknowledgements
Author thanks Barbara Pasquini, Marc Vanderhaeghen, and Lothar Tiator for numerous discussions while working with pion electroproduction amplitudes, Kevin Quirion and Emilie Passemar for numerous discussions while working with pion neutrinoproduction. This work is supported by the US Department of Energy through the Los Alamos National Laboratory and by LANL’s Laboratory Directed Research and Development (LDRD/PRD) program under project numbers 20210968PRD4 and 20240127ER. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the US Department of Energy (Contract No. 89233218CNA000001). FeynCalc [152, 153], LoopTools [154], Mathematica [155], and DataGraph [156] were extremely useful in this work. This work is supported by the National Science Foundation of China under Grants No. 12347105, No. 12447101.